### v1

# Do You Get the Hint? Benchmarking LLMs on the Board Game Concept

Abstract
Large language models (LLMs) have achieved striking successes on many benchmarks, yet recent studies continue to expose fundamental weaknesses. In particular, tasks that require abstract reasoning remain challenging, often because they use representations such as grids, symbols, or visual patterns that differ from the natural language data LLMs are trained on. In this paper, we introduce Concept, a simple word-guessing board game, as a benchmark for probing abductive reasoning in a representation that is much closer to LLM pre-training data: natural language. Our results show that this game, easily solved by humans (with a success rate of over 90%), is still very challenging for state-of-the-art LLMs (no model exceeds 40% success rate). Specifically, we observe that LLMs struggle with interpreting other players’ strategic intents, and with correcting initial hypotheses given sequential information updates. In addition, we extend the evaluation across multiple languages, and find that the LLM performance drops further in lower-resource languages (Dutch, French, and Spanish) compared to English.
Do You Get the Hint? Benchmarking LLMs on the Board Game Concept
Ine Gevers  and Walter Daelemans CLiPS, University of Antwerp ine.gevers@uantwerpen.be,  walter.daelemans@uantwerpen.be
1 Introduction
While Large Language models have propagated impressive progress in NLP recently, it remains a point of debate how much knowledge they possess about the world, and how (or if at all) they reason about this knowledge (see for example Chollet et al. (2025); Kaesberg et al. (2025); Ngo et al. (2025); Benchekroun et al. (2023); Dentella et al. (2024)). A creative way to assess such capabilities is to have LLMs play (natural language) games. In this paper, we add to this line of research by introducing the boardgame ‘Concept’ as a task to measure LLMs’ abstract and abductive reasoning skills. Similar to a single-modal game of Pictionary, there are two roles in Concept (see Figure 1 for an example):
Player 1, the ‘clue-giver’, chooses a concept (a word or a short phrase) and uses a predefined set of clues to express this concept to Player 2. Player 1 has to come up with creative ways to combine these constrained clues so Player 2 can reconstruct the concept in a meaningful way.
Player 2, the ‘guesser’, attempts to guess the concept based on the clues that are given. In a game, there can be multiple guessers.
Figure 1: A simplified example of a round of Concept, comparing guesses made by humans and an LLM.
Both players have a shared goal, namely for Player 2 to guess the correct concept. Both Player 1 and Player 2 have to strategically navigate the shared conceptual space, created by the constrained vocabulary introduced by the predefined set of clues. Making use of this predefined set, the clue-giver has the freedom to express the concept however they see fit. Therefore, the process of clue-giving is inherently more imprecise, complicating a fair evaluation of LLMs (i.e., does the model not know the concept, or is it unable to select and combine seemingly unrelated clues in a meaningful way?). Therefore, in our setup, we implement LLMs solely as the guesser, using human-generated clues as input.
We collect game logs of finished games in English, French, Spanish, and Dutch from the online board game platform Board Game Arena [^1] https://boardgamearena.com/gamepanel?game=concept. For each language, we collect 100 games, which amounts to circa 1,000 concepts played per language. This task aims to answer the following research questions:
  1. 1.
    Compared to humans, how well can LLMs infer concepts from natural language clues, selected by a human from a constrained vocabulary? To what extent can LLMs interpret information expressed through linguistic formulations that differ from how such concepts are typically represented in their training data?
  2. 2.
    Are some categories of concepts (e.g., animals, objects, quotes) easier for models?
  3. 3.
    Is the performance on Concept language-dependent? How well do models transfer abstractions to lower-resource languages?
We hypothesize to observe the following:
  1. 1.
    We expect that model performance does not match, but correlates with human performance. Additionally, we expect that reasoning models outperform non-reasoning models on this task.
  2. 2.
    LLMs perform better on everyday and concrete concepts (e.g., objects, animals) compared to infrequent or abstract concepts (e.g., quotes, culturally-dependent concepts).
  3. 3.
    Performance on low-resource languages is worse.
Our results show that, similar to observations made on visual and spatial reasoning tasks such as ARC and SPaRC, LLMs struggle on a simple language game that humans solve easily (94% solve rate for non-expert humans). Specifically, across all tested models, we observe difficulties with (i) interpreting other players’ strategic intent from the clues they provide, and (ii) correcting initial hypotheses when new information becomes available.
The remainder of the paper is structured as follows. In Section 2, we discuss prior literature. Section 3 introduces the methodology regarding data collection, model prompting, and evaluation. In Section 4, we discuss our results on the Concept benchmark [^2] The data is publicly available for research purposes only: https://huggingface.co/datasets/IneG/concept., and Section 5 concludes this study, summarizing the findings and discussing future works.
2 Related Work
The core challenge that is addressed in this work relates to logical inference and abductive reasoning. Historically, this task was approached by creating symbolic knowledge representations for AI, but has ultimately failed to generalize because of the brittleness and incompleteness of knowledge bases. Later, with the rise of representation learning, distributed vector embeddings became the dominant method for encoding knowledge, enabling models to capture semantic relationships in large text corpora (Mikolov et al., 2013). Building on this foundation, the rise of large autoregressive pre-trained transformer models has made it possible to leverage natural language itself as the medium of knowledge representation (Yang et al., 2023). This evolution has opened the way for creative probes to test reasoning skills in LLMs.
Game settings offer interesting perspectives to test AI models’ capabilities, and understanding of a (physical) world. The rise of powerful auto-regressive Large Language Models (LLMs) gave a significant boost to this research direction. LLMs can embody various roles in games: they can be players, non-player characters, assistants, Game Masters, design games, or analyze gameplay data, and so on (Gallotta et al., 2024). (Natural language) games have been used to evaluate core skills of LLMs, in addition to the standardized benchmarking practices in NLP. For instance, situated language understanding is evaluated through dialogue games (Chalamalasetti et al., 2023), theory of mind is tested through collaborative games (Li et al., 2023a), physical common sense is probed by asking models to organize rooms and place objects in appropriate locations (Zhuo and Murata, 2024) and world knowledge can be tested through the popular game ‘Twenty Questions‘ (De Bruyn et al., 2022). It has also been shown that LLMs as players do not necessarily match human performance, but their performance correlates strongly with difficulty as indicated by humans (Xiao and Yang, 2024).
An interesting subsection of games that are implemented for LLMs are games in which there is a perfect common interest between players. In this setup, there are usually two players: the ‘sender’, and ‘receiver’. The sender disposes of a fixed set of signals. The receiver can observe the signals that are sent, but not the state of the world. A message is encoded in signals by the sender, and the receiver must then try to decode them and decide on which actions to take. For each state, there is only one correct action and both the sender and the receiver have the same outcome objective, namely for the receiver to perform the correct action. Prior research has presented such games as a method to evaluate automated methods, since they test a system’s ability to strategically encode and/or decode information, align meaning between agents, and coordinate actions within a shared framework.
For instance, LLMs have made considerable gains in solving crossword puzzles, especially with the implementation of search algorithms (Saha et al., 2024). However, cryptic puzzles (e.g., relying on wordplay) are much more challenging, and there remains a large performance gap compared to humans (Saha et al., 2024; Sadallah et al., 2024). Leng et al. (2025) create a text+grid crossword benchmark, and show that reasoning models outperform non-reasoning models. However, they do not mention a human baseline. Similarly, the game ‘Codenames’ has been introduced as benchmark to test language nuances and theory of mind, on which LLMs again outperform older methods. However, Stephenson et al. (2025) note that smaller models do not adhere to the constraints of game rules, and models struggle overall with lateral thinking. On the popular puzzle ‘Wordle’, LLMs again underperform but CoT-prompting increases performance (Xiao and Yang, 2024).
However, these games are based on static situations. Lately, interactive games have been used to gauge models’ continuous understanding updates, context-aware responses, and the inference of hidden states of other agents (Momentè et al., 2025). It is argued that these games offer a better proxy for the real-world necessity of adaptability and robustness in interactions. Such game-based evaluations have demonstrated higher correlations with a range of cognitive abilities, including causal and logical inference, as well as social and emotional competencies (Momentè et al., 2025). For instance, in ‘Guess The Build’, players construct a Minecraft build and others try to guess it with natural language. Results show a considerable performance gap between open and API Visual Language Models (VLMs). The performance correlates with frequency of the concept in the training data, and underrepresented cultural concepts and low-resource languages remain an open problem (Zhu et al., 2025). Similarly, the game ‘Pictionary’ has been adapted as a multi-modal benchmark ‘Iconary’, in which a guesser tries to identify a concept that is drawn by composing a set of icons, and the drawer can update their drawing to help the guesser. Results show that T5-models are better at guessing compared to drawing, and the main challenge were out-of-vocabulary words (Clark et al., 2021). Pictionary has also been played by VLMs using the dataset and agentic framework ‘Sketchtopia’, where the win rate is lower compared to humans, and the performance is again linked to frequency in training data (Khan and Sarvadevabhatla, 2025).
In order to be successful in such games, being able to estimate mental states of other players (such as their knowledge, beliefs, or desires), known as Theory of Mind (ToM), is important. While Kosinski (2024) observes that these skills are improving with newer and larger LLMs, Akata et al. (2025) find that LLMs often fail in game settings where cooperation with other players is required, suggesting that ToM understanding of other players is limited. Similarly, Li et al. (2023b) find long contexts and hallucinations are barriers to ToM in LLMs for multi-agent collaboration in text games. Agashe et al. (2023) show that LLMs "perform well in multi-agent coordination settings where decision-making primarily relies on environmental variables but face challenges in scenarios requiring active consideration of partners’ beliefs and intentions".
We contribute to this line of research by introducing ‘Concept’ as a benchmark to evaluate abductive reasoning in LLMs. In this game, the sender selects clues to maximize the probability of the receiver correctly identifying the concept, while constrained to the set of available clues. The receiver, in turn, strategically interprets these symbolic clues, inferring the most plausible concept that best satisfies all of them. The effectiveness of the clues relies on how well the sender’s intended meaning aligns with the receiver’s interpretation. The players must converge on a shared understanding of the concept based on the limited, symbolic clues. This highlights the importance of not just linguistic comprehension but also the strategic intent behind clue selection and interpretation. In order to perform well in Concept, it is not enough for AI to understand each clue separately, but it must be able to meaningfully combine these clues, within the shared conceptual framework that humans use.
‘Concept’ has the advantage of being a very straightforward task, for which there is no need for expert knowledge. Additionally, our setup allows us to compare performance with a human baseline, and the inclusion of multiple languages offers insights into the transferability of LLM’s abduction capabilities. In this way, our benchmark complements popular reasoning challenges such as ARC (Chollet et al., 2025) and SPaRC (Kaesberg et al., 2025). The ARC (Abstraction and Reasoning Corpus) benchmark evaluates generalization through visual pattern-recognition puzzles that require discovering abstract rules from small sets of examples, while SPaRC (Spatial Pathfinding Reasoning Challenge) builds on this idea with text-based grid environments that demand spatial reasoning and step-by-step planning to navigate toward a goal. Both benchmarks probe general reasoning skills but rely on visual or spatial representations that differ substantially from the linguistic data on which LLMs are trained. In contrast, because Concept relies solely on natural language, its representation is more closely aligned with the pre-training data of LLMs, thereby providing a fairer test of their reasoning capabilities.
3 Methodology
3.1 Data
We collected game logs of Concept from the online boardgame platform Board Game Arena (BGA) [^3] https://en.boardgamearena.com/gamepanel?game=concept. Only completed games were kept, and all player data was anonymized (i.e., we removed player ID, timing, stats, etc.). For each language in our dataset (English, French, Dutch, and Spanish), we collected 100 game logs, which amounts to circa 1,000 concepts played per language. In this section, we will detail the structure of our dataset.
Each game includes multiple rounds, with one concept per round. On average, one game consists of 12 concepts. While each concept within a single game is unique, the same concept can appear across different games; in our dataset, the most frequent concept occurs nine times within one language. Importantly, even when the target concept is repeated, the clues differ, since they are provided by different players in different games.
Player 1’s clues are structured hierarchically. The opening clue establishes the top-level category (e.g., man, object, quote), and subsequent clues may either refine this category or form nested sub-hierarchies. In the game interface, these hierarchies are visually marked with colors: each color contains a distinct high-level clue (a pawn), with its own set of subordinate clues (cubes). Player 1 can choose to add extra clues, or remove clues that were given earlier from the board. For each round, we collected (i) the concept; (ii) the clues that were given by Player 1 (with hierarchical structure indicated by color); (iii) the guesses made by Player 2.
On the collected dataset, we filter out the rounds for which there were no clues given (i.e., Player 2 guessed the concept before Player 1 gave any clues). In addition, some rounds contain very few clues because Player 1 removed most of them at the end of the game. To account for this, we report results on two versions of the dataset: one containing all rounds, and one excluding incomplete rounds. Our filtering heuristic considers a round incomplete if the initial clue (the top-level and most important clue) is missing at the end of the game. Table 1 provides an overview of the resulting dataset sizes.
excluded rounds total rounds filtered rounds
English 72 1,103 818
French 89 1,036 829
Dutch 40 1,155 932
Spanish 28 924 716
Table 1: Dataset sizes: excluded rounds (no clues given); total number of rounds (rounds remaining after excluding empty clues); filtered rounds (only complete rounds, in which the first clue is still present at the end of the game).
There is a large variety in how many guesses humans made to either find the correct word or give up, as can be seen in Figure 2. The majority of the words only needed <10 guesses.
Figure 2: Distribution of number of guesses by humans.
3.2 Prompting
We include a wide range of LLMs, both open and closed-sourced: LlaMA3.3 70B; LlaMA4-Maverick; Qwen2.5 72B; Deepseek V3; Mistral Small 24B; and GPT 4.1 mini. Additionally, we include the GPT-OSS 120B reasoning model. These models were chosen because they are among the most recent high-performing releases across major model families, spanning diverse architectures and parameter scales, offering a representative overview of current LLM capabilities. In Appendix A, we include more model details.
For non-reasoning models, we allow 10 output tokens, and keep the temperature stable at 0.1 to keep the output more deterministic and reproducible. For GPT-OSS-120B, we allow 1,000 output tokens, with medium reasoning effort, and follow the recommended temperature setting (t=1). Because of hardware limitations, we prompt the models using the Together.ai API [^4] https://www.together.ai/, except for GPT 4.1 mini, for which we use the OpenAI API [^5] https://platform.openai.com/docs/models/gpt-4.1-mini. The total cost for prompting the models was 120\$. We keep the prompt identical for all models. We organize the clues from Player 1 in a hierarchical structure (grouping clues of the same color together), since tests on the separate validation set (n=50) showed that this technique results in better performance. There is no limit to the number of clues that can be given, as is the case in the original human game. We also include one example in the prompt, taken from the rule book of the original game. Prior works found that one example can work better than multiple examples, ‘possibly due to cognitive overload or excessive focus on example analysis instead of task solving’ (Kaesberg et al., 2025; Ye et al., 2023).
We implement two prompts for the main experiments (in English). Specifically, we create a static and a dynamic prompt. The static prompt consists of the final set of clues given by Player 1 (similar to a snapshot of the clues as they were at the end of the round), and any incorrect guesses made by the LLM are added. We allow 10 guesses per round. This informs us on how well LLMs can navigate the conceptual space, created by humans, and how well they can combine symbolic clues to infer the correct answer. However, as indicated by prior research, including the interaction between players gives more insights into Theory of Mind capabilities. Specifically, in Concept, Player 2 should keep in mind the strategic intent of Player 1’s choice of clues, and why Player 1 might choose to update their clues given Player 2’s incorrect answers. Therefore, in the dynamic prompt, we prompt iteratively, giving the clues from Player 1 per step (one step = the clues given until the original (human) Player 2 interrupted by making a guess). Then, after each step, we add the incorrect guesses from the human players as well as the LLM’s, and update the clues. In this setting, we allow as many guesses as there were made by humans in the original game. We include the prompts in Appendix B.
For the multilingual prompting, we translated the English prompt to the target language [^6] Each translation was proofread by at least a CEFR B2 (ACTFL Advanced Mid) speaker.. We compared translating the entire prompt to the target language, and keeping the instructions in English on a validation set (Vatsal et al., 2025). Given that the latter yielded better results, on the test set we prompt with the entire prompt in the target language. We include the prompts in Appendix B.
3.3 Evaluation
We evaluate model performance using precision at k (prec@k), where a prediction is considered correct only if the guessed concept exactly matches the target concept. To ensure fair comparison, we apply extensive normalization to the LLM outputs, including correcting misrepresented characters (e.g., ASCII artifacts in non-English concepts), and removing formatting artifacts such as the prefix ‘**Concept**:‘ when present. For the static prompt setting, we fix k at 10. For the dynamic prompt setting, k is set to the number of guesses made by humans in the corresponding round (see Figure 2).
4 Results
It is clear that Concept is an easy task for humans. To some extent, for persons who play the game very frequently, there might be a learning effect, but since there are 990 unique concepts per language in the game and our dataset includes 100 unique game plays, we can assume this effect is negligible. Considering all rounds, humans achieve a 92% found rate, and 93% on the filtered rounds. 63% of the rounds were solved within 10 guesses, although the correct answer was often still reached when more guesses were required. As expected, there was a negative correlation (Pearson’s r=0.3r=-0.3) between the number of clues provided and the likelihood of success, which could partially be explained because more difficult concepts requiring more clues. To examine whether performance varied across types of concepts, we categorized rounds by their first (and most important) clue (e.g., wildlife/animal, object/thing/package, building/construction/city, music/song/note), focusing on the top 20 most frequent categories. A Bayesian mixed-effects logistic regression model (GLMM) [^7] https://bambinos.github.io/bambi/notebooks/getting_started.html#generalized-linear-mixed-models was then fit with ‘found’ (i.e., whether the concept was found) as the outcome and category as a predictor [^8] Posterior estimates were obtained with NUTS sampling (4 chains; 1,000 warmup and 1,000 post-warmup draws), and category effects were considered meaningful when their 95% highest-density intervals excluded 0.. The analysis revealed no meaningful differences across categories, indicating that human performance was consistently high regardless of concept type.
In comparison, all LLMs perform significantly worse than humans across the board (see Table 2). Interestingly, in rounds where both humans and the LLM found the correct concept, the LLM overall needs fewer guesses to do so, suggesting greater efficiency once they are on the right track (see Figure 3).
Figure 3: Comparison of how many guesses were needed to arrive at the correct concept, for concepts that are found both by humans and the LLM.
However, in rounds where the correct answer is not found easily, LLMs often get stuck pursuing wrong directions, whereas humans are better at revising their earlier interpretations and still arriving at the correct answer [^9] For models, we observe a strong negative correlation between found rate and number of guesses (Pearson’s $r$ between -0.68 and -0.74), whereas for humans the correlation is weak (r=0.14r=-0.14).. This finding echoes Chollet’s 2019 view of exploration as a key component of intelligence: humans can flexibly revise their hypotheses, but LLMs tend to remain fixated. Since Concept requires interpreting strategic linguistic signals within a shared conceptual space, success depends not only on linguistic fluency but also on adaptive exploration—an area where humans still hold a clear advantage.
To gauge how models compare to each other and to humans on which concepts they (do not) find, we calculate Pearson’s R on the found rates, based on the static prompt. We find that GPT-OSS-120, the reasoning model, has the strongest correlation with human found rates ($r=0.14$). Especially LlaMA4-Maverick has a low correlation with human performance ($r=0.05$). Interestingly, comparing model performance to each other, the correlations are very low, suggesting that different models succeed on largely different subsets of concepts.
Contrary to our expectations, the dynamic prompt yields worse results compared to the static prompt, even though this prompt entails more information, and provides the added possibility to derive useful information from the interaction between Player 1 and Player 2. Corroborating prior research, this suggests that LLMs struggle with simulating Player 1’s strategic intents when giving the clues in in-context learning. Further, as mentioned earlier, even with more explicit representations of the updates in clues and incorrect guesses, LLMs are seemingly unable to update their interpretations of the initial clues, and remain stuck in the wrong direction.
Both for the static and dynamic prompt, we observe instances where the model repeats its previous (incorrect) answer, even though the prompt is updated with this information. Additionally, there are rounds where the guesses are not valid given the constraints of the highest level clue: for instance, if the first clue is ‘building’, the LLM’s guesses are not buildings.
Again contrary to our initial expectations, the reasoning model we included in our experiments (GPT-OSS-120B) does not outperform non-reasoning models. In some cases, this is caused because the number of reasoning tokens (n=1,000) were insufficient and the model did not arrive at an answer. This finding agrees with previous results that longer reasoning processes, in which the model generates more intermediate tokens, does not necessarily help the model to arrive at useful associations or correcting the erroneous directions (Yuan et al., 2025; Hassid et al., 2025; Vamvourellis and Mehta, 2025).
Table 3 shows the comparison of human and LLM performance per category (grouped by the first, highest-level clue). To compare LLMs with humans across categories, we extended the GLMM model to include solver (human vs. LLM), high-level-clue category, and their interaction as fixed effects. This allowed us to test whether differences in success rates between humans and LLMs varied systematically by concept category. For most models, there is no significant difference, except for Qwen2 72B (worse on ‘expression/quote, speak, word’, and ‘theater/film/camera’), LlaMA4-Maverick (worse on ‘expression/quote, speak, word’), and Mistral-Small 24B (worse on ‘expression/quote, speak, word’, and ‘object/thing/package’). To some extent, this agrees with our initial hypothesis that LLMs would struggle more with infrequent and culturally-dependent concepts, such as expressions and quotes.
Finally, we observe similar trends for Dutch, French, and Spanish, although the differences are even more pronounced. While human performance remains consistently high (94% found rate), all LLMs perform worse, and the gap relative to humans is larger than in English (see Table 4). Between the non-English languages, models seem to perform better on Spanish. This could be a reflection of common internet language distributions, where Spanish has a larger share compared to the other two languages. However, the models’ documentation does not provide detailed frequency distributions for non-English training data. Most model specifications list French and Spanish as supported languages, but not Dutch, yet LLMs’ performance on Dutch is not significantly worse than on French. This is particularly surprising for Mistral, given its French origins. Overall, the weak performance of LLMs on non-English Concept games highlights their limited capacity to transfer knowledge for this type of abstraction.
Human llama3-70B Qwen2.5-72B Deepseek-V3 llama4-maverick Mistral-24B GPT-4.1 mini GPT-OSS-120B
Aggregated (static) 0.9184 0.3469 0.2982 0.3463 0.2747 0.2955 0.3445 0.2792
Aggregated (dynamic) 0.9184 0.2533 0.2336 0.3054 0.2246 0.2075 0.2893 0.2318
Only complete (static) 0.9339 0.3924 0.3496 0.3863 0.3117 0.3337 0.3728 0.3129
Only complete (dynamic) 0.9339 0.2713 0.2603 0.3190 0.2432 0.2139 0.3068 0.2493
Table 2: Found rates per model on the English dataset. Entire dataset: n=1,103, filtered: n=818.
Human llama3-70B Qwen2.5-72B Deepseek-V3 llama4-maverick Mistral-24B GPT-4.1 mini GPT-OSS-120B
Wildlife, animal (n=83) 0.9759 0.5662 0.5903 0.5421 0.4337 0.5542 0.5180 0.3855
Theater, film, camera (n=78) 0.9358 0.3974 0.1282 0.2948 0.2051 0.2307 0.2564 0.1794
Expression/quote, speak, word (n=69) 0.8550 0.1594 0.0724 0.2318 0.0579 0.1014 0.1304 0.2028
Object, thing, package (n=57) 0.9824 0.4561 0.4210 0.3684 0.3508 0.2280 0.4736 0.3333
Food, nutrition, edible (n=47) 0.9361 0.4893 0.4042 0.4468 0.3829 0.4893 0.5319 0.4893
Male/man, husband, masculine (n=43) 0.9302 0.2558 0.2790 0.2558 0.2093 0.3023 0.2790 0.3023
Fictional, imaginary, wish (n=28) 0.8214 0.3571 0.3928 0.3928 0.3928 0.3928 0.5357 0.3571
Building, construction, city (n=27) 0.9259 0.3333 0.4814 0.4814 0.2592 0.3703 0.3333 0.2222
Literature, writing, book (n=24) 0.9583 0.3333 0.2916 0.3333 0.2916 0.2916 0.4583 0.2500
Television, broadcast, series (n=22) 0.9545 0.3181 0.2727 0.4545 0.3181 0.2727 0.3181 0.4090
Person, family, group (n=20) 0.9500 0.3000 0.1500 0.2500 0.2000 0.1000 0.2000 0.3000
Female/woman, wife, feminine (n=18) 0.9444 0.3888 0.2777 0.3888 0.2777 0.3333 0.3333 0.3888
Games, toys, youth (n=16) 0.8750 0.5000 0.4375 0.6250 0.5625 0.3750 0.56250 0.5000
Flora, plant, nature (n=15) 1.000 0.4000 0.3333 0.2666 0.3333 0.4000 0.4000 0.3333
Location, country, flag (n=13) 0.9230 0.3846 0.4615 0.4615 0.3846 0.3076 0.3076 0.3846
Water, liquid, aquatic (n=11) 0.9090 0.5454 0.5454 0.5454 0.5454 0.4545 0.4545 0.4545
Music, song, note (n=11) 0.9090 0.0909 0.0909 0.0909 0.0909 0.0000 0.0909 0.1818
Work, profession, craft (n=11) 1.000 0.4545 0.6363 0.4545 0.5454 0.3636 0.1818 0.0909
Tools, construction (n=11) 1.000 0.5454 0.4545 0.7272 0.3636 0.7272 0.8181 0.0909
Table 3: Found rates per category (for each category with over 10 observations) for each model. We report this for the static prompt, on the filtered dataset (n=818) since this is the best setting across all models.
Language Split Human llama3-70B Qwen2.5-72B Deepseek-V3 llama4-maverick Mistral-24B GPT-4.1 mini GPT-OSS-120B
French Aggregated (static) 0.9403 0.1462 0.1583 0.2123 0.2123 0.1438 0.2133 0.2133
Only complete (static) 0.9457 0.1712 0.1845 0.2460 0.2448 0.1640 0.2376 0.2279
Dutch Aggregated (static) 0.9402 0.2008 0.1696 0.2320 0.2112 0.1350 0.2363 0.2363
Only complete (static) 0.9484 0.2167 0.1856 0.2371 0.2210 0.1480 0.2435 0.2414
Spanish Aggregated (static) 0.9469 0.2208 0.2218 0.3131 0.2622 0.2239 0.2823 0.2845
Only complete (static) 0.9497 0.2555 0.2486 0.3351 0.2877 0.2513 0.3100 0.3114
Table 4: Found rates per model for French, Dutch, and Spanish. Since the static prompt consistently outperforms the dynamic one, we only include results for the static prompt. Entire dataset French: n=1,036, filtered: n=829. Entire dataset Dutch: n=1,155, filtered: n=932. Entire dataset Spanish: n=942, filtered: n=716.
5 Conclusion
In this paper, we present the game ‘Concept’ as a test for abductive reasoning. This game is comparable to Pictionary, but relying solely on natural language: Player 1 selects clues from a predefined set to describe a concept, Player 2 has to guess this concept. Whereas previous reasoning challenges such as ARC and SPaRC have addressed shortcomings in LLMs’ reasoning capacities, Concept levels the playing field by having the task representations rely only on natural language, which is better aligned with the pre-training data of LLMs.
We collected game logs in English, Dutch, French, and Spanish from the online boardgame platform Board Game Arena, providing a valuable human baseline performance. Per language, this results in about 1,000 concepts. We released the data for research purposes.
In our experiments, we prompted 7 state-of-the-art LLMs, both open- and closed-source. We implemented two prompting strategies, static and dynamic. The former presents the final set of clues given by Player 1, while the latter provides a level of interaction by presenting the clues iteratively, as they were given in the original game.
This language game, which is very easy for humans (with success rates over 90%), reveals important shortcomings of current LLMs. Specifically, we confirm previous findings that LLMs struggle to interpret Player 1’s strategic intents when giving the clues, which is supported by the finding that the dynamic prompt does not result in better results. Further, LLMs struggle to update their initial interpretations of the clues when sequential information updates are given. Even though incorrect guesses made by the model are added to the prompt, this does not help the model to arrive at the correct answer: often, the model remains stuck in the wrong direction. Additionally, we noted that in cases where both the LLM and humans identified the concept, LLMs consistently require fewer guesses, but in cases where more guesses are needed, humans are better at correcting their abstractions to still find the correct answer.
Comparing results from the dominant English language to lower-resource languages (Spanish, French, and Dutch), we observed that LLMs perform worse on non-English languages.
For further research, we suggest fine-tuning an LLM for this task, because previous research has shown that fine-tuning a small set of parameters can drastically increase performance on abstract reasoning tasks such as ACRE and RAVEN (Yun et al., 2025). Further, for interpretability, we suggest investigating the link between found rate of a concept and its frequency in the LLM’s pre-training data, or singling out the importance of one clue by implementing an ablation study. Additionally, since Concept can be played with multiple people guessing, it would be interesting to implement this as a multi-agent setting, in which multiple models guess, and each model sees the output of the others. Given the low correlation between models’ found rates, the interaction between different models could potentially aid performance.
Limitations
This study is subject to a few limitations. First of all, our evaluation metric, Exact Match, might miss more fine-grained insights. For instance, synonyms of the correct answer are now marked as incorrect. Nonetheless, this conservative measure provides a clear, reproducible benchmark. Moreover, even under this strict criterion, humans still achieve a success rate of over 90%, underscoring that the task remains feasible. Second, we observed instances where the LLM did not adhere to the instructions of only outputting the final answer, and is interrupted mid-stream due to the limited number of allowed output-tokens. Since the model was instructed to return only the answer as output, we rule this behavior a task failure, and count these cases as incorrect. Third, prompting various models inevitably introduces a trade-off between reproducibility and model sensitivity. We chose to maintain an identical prompt across all models to ensure stability and reproducibility, but we acknowledge that different models might have performed better with variations of the prompt.
Acknowledgments
This research was made possible with a grant from the Fonds Wetenschappelijk Onderzoek (FWO) project 42/FA030100/9770. We would like to thank Victor De Marez for his practical support in retrieving the game logs from the Board Game Arena platform. Also many thanks to Pieter Fivez and Luna De Bruyne for insightful brainstorming and discussions.
Appendix A Model specifications
In Table 5, we detail model name, version, and release date of the models included in our experiments.
Model Full Name / Variant Size Release Date Source
LLaMA-4 Maverick LLaMA-4-Maverick-17B-128E-Instruct-FP8 17B active (400B total) Apr 2025 https://api.together.ai/models/meta-llama/Llama-4-Maverick-17B-128E-Instruct-FP8
LLaMA-3.3 LLaMA-3.3-70B-Instruct-Turbo 70B Dec 2024 https://api.together.ai/models/meta-llama/Llama-3.3-70B-Instruct-Turbo
Qwen-2.5 Qwen-2.5-72B-Instruct-Turbo 72B Sept 2024 https://api.together.ai/models/Qwen/Qwen2.5-72B-Instruct-Turbo
DeepSeek V3 DeepSeek-V3 671B total, 37B active Dec 2024 https://api.together.ai/models/deepseek-ai/DeepSeek-V3
Mistral Small Mistral-Small-24B-Instruct-2501 24B Jan 2025 https://api.together.ai/models/mistralai/Mistral-Small-24B-Instruct-2501
GPT-4.1 mini gpt-4.1-mini-2025-04-14 n/a Apr 2025 https://platform.openai.com/docs/models/gpt-4.1-mini
GPT-OSS 120B GPT-OSS-120B (Reasoning) 120B Aug 2025 https://api.together.ai/models/openai/gpt-oss-120b
Table 5: Models evaluated in our experiments, with full names, sizes, release dates, and sources. Sizes are as reported in model documentation (for MoE models, both total and active parameters are listed).
Appendix B Prompt
Below, we include the template for the static prompt in English:
[INST]Let’s play Concept! I will give you clues, you have to guess the concept. Only output the concept, nothing else! The hints are organized by color. Per color, the hierarchy of the clues is indicated by ’High-level’ and ’Low-level’. Low-level clues are given in function of the high-level clue of the same color. I will also give you the incorrect guesses, so you can learn from them. Here is an example: **Hints**: We are looking for a concept in this category: Building, Construction, City (highest-level hint, color: green). An additional hint in color green is: Metal. A high-level hint in color red is: Location, Country, Flag. An additional hint in color red is: Blue. An additional hint in color red is: White. An additional hint in color red is: Red. **Concept**: Eiffel Tower
Here are the clues: **Hints** [insert clues here] **Wrong guesses** [insert previous wrong answers here] **Concept**[/INST]
Here, we include the template for the dynamic prompt in English. The main difference with the static prompt is that only clues until a certain step are given, instead of the clues as they appear at the end of a round. Specifically, we define one step as the clues given by Player 1 until Player 2 interrupts by making a guess. For the next step, we update the clues as Player 1 did (adding or removing clues) until Player 2 interrupted again, and so forth. Second, we include both the incorrect guesses made by humans, as well as LLMs for that step.
[INST]Let’s play Concept! I will give you clues, you have to guess the concept. Only output the concept, nothing else! The hints are organized by color. Per color, the hierarchy of the clues is indicated by ’High-level’ and ’Low-level’. Low-level clues are given in function of the high-level clue of the same color. I will also give you the incorrect guesses, so you can learn from them. Here is an example: **Hints**: We are looking for a concept in this category: Building, Construction, City (highest-level hint, color: green). An additional hint in color green is: Metal. A high-level hint in color red is: Location, Country, Flag. An additional hint in color red is: Blue. An additional hint in color red is: White. An additional hint in color red is: Red. **Concept**: Eiffel Tower
Here are the clues: **Hints** [insert clues until the current step here] **Wrong guesses** [insert previous wrong answers from LLMs and humans until the previous step here] **Concept**[/INST]
The static prompt in Spanish:
[INST]¡Juguemos a Concept! Te daré pistas en español y tú tendrás que adivinar el concepto en español. ¡Solo da el concepto, nada más! Las pistas están clasificadas por colores. Para cada color, la jerarquía de las pistas se indica con ’Nivel alto’ o ’adicional’. Las pistas adicionales se dan en función de las pistas de alto nivel del mismo color. También te daré las respuestas incorrectas para que puedas aprender de ellas. Aquí tienes un ejemplo: *Pistas**: Buscamos un concepto en esta categoría: Edificio, Construcción, Ciudad (pista de nivel superior, color: verde). Una pista adicional en color verde es: Metal. Una pista de alto nivel en color rojo es: Ubicación, País, Bandera. Una pista adicional en color rojo es: Azul. Una pista adicional en color rojo es: Blanco. Una pista adicional en color rojo es: Rojo. **Concepto**: Torre Eiffel.
Estas son las pistas: **Pistas** [insert clues here] **Respuestas incorrectas**: [insert previous wrong answers here] **Concepto**[/INST]
The static prompt in French:
[INST]Jouons à Concept! Je vais vous donner des indices en français, vous devez deviner le concept en français. Ne donnez que le concept, rien d’autre ! Les indices sont classés par couleur. Pour chaque couleur, la hiérarchie des indices est indiquée par « Haut niveau » et « Supplémentaire ». Les indices supplémentaires sont donnés en fonction de l’indice de haut niveau de la même couleur. Je vous donnerai également les réponses incorrectes, afin que vous puissiez en tirer des leçons. Voici un exemple: **Indices**: Nous recherchons un concept dans cette catégorie: Bâtiment, Construction, Ville (indice de plus haut niveau, couleur : vert). Un indice supplémentaire de couleur vert est: Métal. Un indice de haut niveau de couleur rouge est: Lieu, Pays, Drapeau. Un indice supplémentaire de couleur rouge est: Bleu. Un indice supplémentaire de couleur rouge est: Blanc. Un indice supplémentaire de couleur rouge est: Rouge. **Concept**: Tour Eiffel
Voici les indices: **Indices** [insert clues here] Réponses incorrectes**: [insert previous wrong answers here] **Concept**[/INST]
The static prompt in Dutch:
[INST]Laten we Concept spelen! Ik ga je aanwijzingen geven in het Nederlands, je moet het concept in het Nederlands raden. Geef alleen het concept, niets anders! De hints zijn gecategoriseerd op kleur. Voor elke kleur is de hiërarchie van de hints aangegeven door ’Hoog niveau’ of ’Extra’. De extra hints worden gegeven in functie van de hints op hoog-niveau van dezelfde kleur. Ik zal je ook de foute antwoorden geven, zodat je hiervan kunt leren. Hier is een voorbeeld: **Hints**: We zoeken een concept in deze categorie: Gebouw, Bouw, Stad (hoogste niveau hint, kleur: groen). Een extra hint in kleur groen is: Metaal. Een hint op hoog niveau in kleur rood is: Locatie, Land, Vlag. Een extra hint in kleur rood is: Blauw. Een extra hint in kleur rood is: Wit. Een extra hint in kleur rood is: Rood. **Concept**: Eiffeltoren
Hier zijn de hints: **Hints** [insert clues here] **Foute antwoorden**: [insert previous wrong answers here] **Concept**[/INST]